OpenStudy (anonymous):

The maximum height a model rocket reaches that is fired from the ground straight up into the air varies directly with the square of the time it takes for the model rocket to return to the ground. If a model rocket reaches a maximum height of 512 feet, and returns to the ground 8 seconds after it was fired, what is the constant of variation?

OpenStudy (tad1):

h =kt^2 512 = k (8)^2 512 = 64k k = 512/64 k = ? don't have a calculator